Excess dilute HCl was added to sodium sulphite 960cm3 of sulphur IV oxide has was produced calculate the mass of the sulphate used
1
Expert's answer
2021-09-10T00:41:48-0400

Putting molar gas volume ="24000cm^3" and molar mass of sulphite =126g, mass of the sulphite used can be calculated as follows.

number of moles of sulphur (iv) oxide "=\frac{960cm^3 \times 1mole}{24000cm^3}=0.04moles"

Equation:

"Na_2SO_3 (s)+2HCl_{aq} \to 2NaCl_{aq} +SO_2(g)+H_2O_(l)"

mole ration "Na_2SO_3:SO_2" is 1:1

number of moles of "Na_2SO_3=0.04moles"

mass of "Na_2SO_3=126gmol^{-1} \times 0.04=5.04g"